OpenStudy (anonymous):

WILL MEDAL Samantha puts an initial $500 into a savings account. The account has a 4% annual compound interest rate. What is the function that best represents Samantha's account and how much does she have in savings after 5 years? Round your answer to the nearest dollar. Hint: Use the formula, f(x) = P(1 + r)x. f(x) = 500(1.4)x, $2,689 f(x) = 500(0.04)x, $512 f(x) = 500(1.04)x, $608 f(x) = 500(1.04)x, $527

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

hint: f(x) = P(1 + r)x. P = initial amount r = interest rate

OpenStudy (anonymous):

Would it look like this 500 ( 1+4%) = x @Vocaloid

OpenStudy (danjs):

yes, but put 4% into a decimal so you can add it to the 1 each year you multiply your current amount by 104%, (1+r)

OpenStudy (anonymous):

so if its in decimal form it would be 1.04?

OpenStudy (danjs):

right

Vocaloid (vocaloid):

yeah

OpenStudy (danjs):

(1.04)^x each year x increases, multiplies another (1.04) again 500(1.04)^x

OpenStudy (anonymous):

okay so would the answer be c?

OpenStudy (danjs):

c or d, what it is when x=5,

OpenStudy (anonymous):

C right

OpenStudy (danjs):

$608.33

OpenStudy (danjs):

yea

OpenStudy (anonymous):

Sweet I understanding thx a lot.